Suppose that a consumer with income of birr 90 wants to spend the whole income on good B and good A. Assume that the goods are consumed on a one-to-one basis and one good is worthless without the other. The price of good A and B are both 3 birr.
A. How many units of good A&B will the consumer buy?
B. If the price of good A increases to birr 6, how many he/she buy of each?
C.vWhat type of preferences does consumption of the two goods represent? What is the shape of the indifference curve for this type of preference?
1
Expert's answer
2021-08-13T13:39:16-0400

Income M = "90"

Quantities "A \space and \space B"

M = QaPa+ QbPb

"Qa = \frac{M- QbPb}{Pa}"

"Qa = \frac {M}{Pa} - \frac {QbPb}{Pa}" This is the equation of the budget line.


Qa ="\frac {90}{3} - \frac {QbPb}{Pa}"


"Qa = 30 - \frac {3Qb}{3}"

Assuming the 1:1 ratio of usage, 15 units of A and 15 units of B will be bought at the price of 3 birr

"15 = 30 - Qb"

"Qb = 15\space units"

"Qa= 15\space unit"

When the price is increased to 6 birr while maintaining the 1:1 usage of two commodities

1 commodity of A = 6 birr

1 commodity of B = 3 birr

One usage costs= 9

To find the consumed pair of commodities, we compute "\frac{90}{9}= 10"

Therefore the quantities are;

"Qb = 10 \space units"

"Qa = 10 \space units"


The kind of preference is the perfect complements.

This type of preference has the right angle indifference curve as shown below.